# Group-Relative REINFORCE Is Secretly an Off-Policy Algorithm: Demystifying Some Myths About GRPO and Its Friends

### v1

@@ 4.3 Understanding data-weighting methods
The idea is to use a filtered subset $\mathcal{S}\subseteq[K]$ of responses for training; for example, the Kimi-Researcher technical blog (Kimi-Team, 2025a) proposes to “discard some negative samples strategically”, as negative gradients increase the risk of entropy collapse. This is indeed a special case of Eq. (9), by setting $w_{i}=\sqrt{K}/|\mathcal{S}|$ for $i\in\mathcal{S}$ and 0 otherwise:
@@
This is the pairwise-weighted REINFORCE gradient in Eq. (9), plus a regularization term (weighted by $\overline{r}_{w}-\overline{r}>0$) that resembles the one in AsymRE but prioritizes imitating higher-reward responses, echoing the finding from offline RL literature (Hong et al., 2023a, b) that regularizing against high-reward trajectories can be more effective than conservatively imitating all trajectories in the dataset.
Implementation details.
Below are the concrete instantiations adopted in our empirical studies:
  • RED-Weight: Each sample $i$ is weighted by $w_{i}=\exp({A_{i}}/{\tau})$, where $A_{i}$ denotes its advantage estimate and $\tau>0$ is a temperature parameter controlling the sharpness of weighting. This scheme amplifies high-advantage samples while down-weighting low-advantage ones. We fix $\tau=1$ for all experiments.
Figure 5: Empirical performance of RED-Drop and RED-Weight on GSM8k with Qwen2.5-1.5B-Instruct, in both on-policy and off-policy settings. Training reward curves are smoothed with a running-average window of size 3.
@@
Figure 5 presents GSM8k results with Qwen2.5-1.5B-Instruct, which confirm the efficacy of RED-Drop and RED-Weight in on/off-policy settings, comparable to REC-OneSide-NoIS with enlarged $(\epsilon_{\textsf{low}},\epsilon_{\textsf{high}})$. Figure 6 reports larger-scale experiments on Guru-Math with Qwen2.5-7B-Instruct, where RED-Weight achieves higher rewards than GRPO, with similar KL distance to the initial policy. Figure 7 further validates the efficacy of RED-Weight on MATH with Llama-3.1-8B-Instruct; compared to GRPO and REC-OneSide-NoIS, RED-Weight achieves higher rewards with lower KL divergence, while maintaining more stable entropy and response lengths.
Figure 7: Comparison of RED-Weight, REC-OneSide-NoIS, and GRPO on MATH with Llama-3.1-8B-Instruct. Reported metrics for training include reward, KL distance to the initial model, entropy, and response length. We also report evaluation accuracy on the MATH500 subset.
5 Related works
@@
Off-policy RL for LLMs has been studied from various perspectives. Importance sampling has long been considered one foundational mechanism for off-policy RL; besides PPO and GRPO, recent extensions include GSPO (Zheng et al., 2025) and GMPO (Zhao et al., 2025) that work with sequence-wise probability ratios, CISPO (Chen et al., 2025) that clips probability ratios rather than token updates, decoupled PPO (Fu et al., 2025a) that adapts PPO to asynchronous RL, among others. AsymRE (Arnal et al., 2025) offers an alternative baseline-shift approach (with ad-hoc analysis for discrete bandit settings), while OPMD (Kimi-Team, 2025b) partly overlaps with our analysis up to Eq. (4) before diverging, as discussed earlier in Section 4.2. Contrastive Policy Gradient (Flet-Berliac et al., 2024) overlaps with our analysis up to Eq. (6), but it requires paired responses within the same micro-batch (in order to optimize the pairwise surrogate loss), rendering it less infra-friendly than REINFORCE variants. Other perspectives include learning dynamics of DPO and SFT (Ren and Sutherland, 2025), training offline loss functions with negative gradients on on-policy data (Tajwar et al., 2024), or improving generalization of SFT via probability-aware rescaling (Wu et al., 2025). Another line of research integrates expert data into online RL (Yan et al., 2025; Zhang et al., 2025c; Fu et al., 2025b). Our work contributes complementary perspectives to this growing toolkit for off-policy LLM-RL.
6 Limitations and future work
@@
While our work offers a new off-policy interpretation for group-relative REINFORCE and shows its broad implications for LLM-RL, several limitations remain. (1) Our current analysis covers single/multi-step RL with response/trajectory-level rewards, and assumes access to multiple rollouts per query. Future work may expand its scope and applicability, e.g., generalizing to settings with step-level rewards or only one rollout per query. (2) Our analysis lacks formal guarantees for policy improvement or convergence. Future work may identify distributional assumptions that yield provable guarantees for REINFORCE variants in off-policy settings. (3) Our experiments focus on settings where training data is generated by older policy versions. Extensions to broader off-policy settings (e.g., advanced experience synthesis or incorporation of expert data) may reveal new insights. Addressing these limitations will further solidify the theoretical foundation and advance principled algorithm design for off-policy LLM-RL.
@@ Appendix B Implementation details and additional experiments
We implement all algorithms with the Trinity-RFT framework [Pan et al., 2025], and run experiments on NVIDIA L20, H20, and A800 GPUs. See Tables 1 and 2 for detailed configurations of our experiments.
@@ B.1 Dataset details
  • GSM8k [Cobbe et al., 2021] is a widely used benchmark with 8.5k grade-school math word problems, designed to test arithmetic reasoning and step-by-step problem solving.
  • MATH [Hendrycks et al., 2021] covers algebra, geometry, probability, and number theory, containing 12.5k examples in total (7.5k for training and 5k for testing); it demands advanced symbolic reasoning beyond GSM8k.
  • Guru [Cheng et al., 2025] is a multi-domain reasoning dataset with 91.9k examples spanning math, code, science, logic, simulation, and tabular tasks; we use its math subset (around 54k samples), which introduces diverse problem formats for evaluating transfer of reasoning strategies.
  • ToolACE [Liu et al., 2025a] is a multilingual benchmark with around 11k synthetic samples designed to evaluate LLMs’ ability to solve tasks by selecting and invoking external tools via strict JSON-formatted function calls; we use a 5k single-turn subset in our experiments.
@@ B.2 Understanding the synchronization parameters
The parameter sync_interval specifies the number of generated rollout batches (which equals the number of gradient steps for training the policy model) between two consecutive executions of model weight synchronization. When sync_interval =1=1, the rollout and policy models synchronize after each gradient step with one batch of samples, yielding a strictly on-policy process (if we ignore the issue of precision mismatch between rollout and training engines [Yao et al., 2025]). When sync_interval >1>1, sync_interval rollout batches are generated with stale model weights before synchronization, which accelerates the overall RL process through pipeline parallelism but incurs off-policyness.
The parameter sync_offset specifies the lag between the generation and consumption of each batch. More specifically, sync_offset batches are generated and saved to the buffer before training is launched, which is also useful for reducing pipeline bubbles and improving hardware utilization [Noukhovitch et al., 2025]. In some of our experiments, we deliberately set sync_offset to a large value, in order to simulate a scenario where reward signals from the environment are lagged.
@@
GSM8K
(Qwen2.5-1.5B)
ToolACE
(Llama-3.2-3B)
Guru
(Qwen2.5-7B)
MATH
(Llama-3.1-8B)
Learning rate $\times\!10^{-6} 1×1$$\times\!10^{-6}$ $1\!\times\!10^{-6}$ $5\!\times\!10^{-7}$
Batch size 96 96 64 64
$K$ 8 8 16 16
Weight decay 0.01 0.01 0.1 0.1
Warmup steps 0 0 80 40
Eval temperature 1.0 N/A N/A 0.6
Eval top-p 1.0 N/A N/A 1.0
Figures 3, 5, 9, 10, 11 4 6 7
@@ B.3 REC with different clipping mechanisms
We replace the mask $M_{i}^{t}$ in REC-OneSide-IS/NoIS in Eq. (8) with a two-side mask [^3] It turns out that REC-TwoSide-NoIS resembles the sPPO algorithm proposed by Vaswani et al. [2022], though derived with different rationales. :
@@
Two-side clipping imposes weaker regularization than one-side clipping does with the same clipping parameter $(\epsilon_{\textsf{low}},\epsilon_{\textsf{high}})$. This can potentially improve training efficiency, but might also be risky when the probability ratio $\pi_{\bm{\theta}}/\pi_{\textsf{old}}$ goes far off. To compensate for this, we design REC-Ring.
@@ B.5 Experiments for OPMD and AsymRE
Figure 11 presents empirical results for OPMD and AsymRE in various off-policy settings. It is worth noting that, while the analysis and experiments in their original papers [Kimi-Team, 2025b, Arnal et al., 2025] focus on a setting that is effectively the same as our $\texttt{sync\_interval}>1$ setting, our analysis and experiments have also validated their efficacy in $\texttt{sync\_offset}>1$ scenarios.
@@ Appendix C Summary: a unified view of various algorithms
Add regularization loss OPMD $\widehat{L}=-\tfrac{1}{K}\!\sum_{i}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}|x)+\tfrac{\tau}{2K}\!\sum_{i}(\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\textsf{old}}(y_{i}|x))^{2}$
AsymRE $\widehat{L}=-\tfrac{1}{K}\!\sum_{i}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}|x)-\tfrac{\tau}{K}\!\sum_{i}\log\pi_{\bm{\theta}}(y_{i}|x)$
@@
RED-Weight $\bm{g}=\sum_{i}w_{i}(r_{i}-\overline{r})\,\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x),\;\;w_{i}=\exp(A_{i}/\tau)$

### v2

@@ 4.3 Understanding data-weighting methods
The idea is to use a filtered subset $\mathcal{S}\subseteq[K]$ of responses for training (Shang et al., 2025). For example, the Kimi-Researcher blog (Kimi-Team, 2025a) proposes to “discard some negative samples strategically”, as negative gradients increase the risk of entropy collapse. This is indeed a special case of Eq. (9), by setting $w_{i}=\sqrt{K}/|\mathcal{S}|$ for $i\in\mathcal{S}$ and 0 otherwise:
@@
This is the pairwise-weighted REINFORCE gradient in Eq. (9), plus a regularization term (weighted by $\overline{r}_{w}-\overline{r}>0$) that resembles the one in AsymRE but prioritizes imitating higher-reward responses, echoing the finding from offline RL literature (Hong et al., 2023a; b) that regularizing against high-reward trajectories can be more effective than conservatively imitating all trajectories in the dataset.
Figure 4: Empirical performance of RED algorithms on GSM8k with Qwen2.5-1.5B-Instruct, in both on-policy and off-policy settings. Training reward curves are smoothed with a running-average window of size 3. Implementation details about RED-Weight and RED-Drop are provided in Appendix B.6.
@@
Figure 4 presents GSM8k results with Qwen2.5-1.5B-Instruct, which confirm the efficacy of RED-Drop and RED-Weight (details in Appendix B.6) in on/off-policy settings, comparable to REC-OneSide-NoIS with enlarged $(\epsilon_{\textsf{low}},\epsilon_{\textsf{high}})$. Figure 5 reports larger-scale experiments on Guru-Math with Qwen2.5-7B-Instruct, where RED-Weight achieves higher rewards than GRPO, with similar KL distance to the initial policy. Figure 11 in the appendix further validates the efficacy of RED-Weight on MATH with Llama-3.1-8B-Instruct.
5 Discussions
Related works.
@@ 5 Discussions
Off-policy RL for LLMs has been studied from various perspectives. Importance sampling has long been considered one foundational mechanism for off-policy RL; besides TRPO, PPO and GRPO, recent extensions include GSPO (Zheng et al., 2025) and GMPO (Zhao et al., 2025) that work with sequence-wise probability ratios, CISPO (Chen et al., 2025) that clips probability ratios rather than token updates, decoupled PPO (Fu et al., 2025a) that adapts PPO to asynchronous RL, among others (Roux et al., 2025; Zheng et al., 2026; Xi et al., 2026; Wang et al., 2025a). AsymRE (Arnal et al., 2025) offers an alternative baseline-shift approach (with ad-hoc analysis for discrete bandit settings), while OPMD (Kimi-Team, 2025b) partly overlaps with our analysis up to Eq. (4) before diverging, as discussed earlier in Section 4.2. REBEL (Gao et al., 2024) and CoPG (Flet-Berliac et al., 2024) overlap with our analysis up to Eq. (6) before diverging, which will be elaborated in Appendix D. Other perspectives for off-policy LLM-RL include learning dynamics of DPO and SFT (Ren & Sutherland, 2025), training offline loss functions with negative gradients on on-policy data (Tajwar et al., 2024), or improving generalization of SFT via probability-aware rescaling (Wu et al., 2025). Another line of research integrates expert data into online RL (Yan et al., 2025; Zhang et al., 2025c; Fu et al., 2025b). Our work contributes complementary perspectives to this growing toolkit for off-policy LLM-RL. Further discussion on prior works that are most closely related to our main analysis can be found in Appendix D.
Limitations and future work.
@@
While our work offers a new off-policy interpretation for group-relative REINFORCE and shows its broad implications for LLM-RL, several limitations remain. (1) Our current analysis covers single/multi-step RL with response/trajectory-level rewards, and assumes access to multiple rollouts per query. Future work may expand its scope and applicability, e.g., generalizing to settings with step-level rewards or only one rollout per query. (2) Our analysis lacks formal guarantees for policy improvement or convergence. Future work may identify distributional assumptions that yield provable guarantees for REINFORCE variants in off-policy settings. (3) Our experiments focus on settings where training data is generated by older policy versions. Extensions to broader off-policy settings, e.g., advanced experience synthesis (Shi et al., 2026) or incorporation of expert data (Yan et al., 2025; Zhang et al., 2025c), may reveal new insights. Addressing these limitations will further solidify the theoretical foundation and advance principled algorithm design for off-policy LLM-RL.
Acknowledgments
The authors would like to thank the anonymous reviewers and Area Chairs for their constructive feedback that has helped improve this work.
Reproducibility statement
Full implementation details and hyperparameter configurations are documented in Section 4 and Appendix B. We have released our code publicly to facilitate reproducibility.
Ethics statement
All datasets used in this study (e.g., GSM8k, MATH, Guru, ToolACE) are publicly available, and no private or personally identifiable information was collected or used. Our contributions are methodological, focusing on improving the stability and efficiency of RL for LLM post-training. We acknowledge that LLMs may still generate biased or harmful outputs; however, our experiments are restricted to benchmark evaluations and do not involve deployment in real-world systems. We believe that releasing our code and reporting detailed hyperparameter settings will foster reproducibility and responsible advancement in this field.
@@ Appendix B Implementation details and additional experiments
We implement all algorithms with the Trinity-RFT framework (Pan et al., 2025), and run experiments on NVIDIA L20, H20, and A800 GPUs. See Tables 1 and 2 for detailed configurations of our experiments.
@@ B.1 Dataset details
  • GSM8k (Cobbe et al., 2021) is a widely used benchmark with 8.5k grade-school math word problems, designed to test arithmetic reasoning and step-by-step problem solving.
  • MATH (Hendrycks et al., 2021) covers algebra, geometry, probability, and number theory, containing 12.5k examples in total (7.5k for training and 5k for testing); it demands advanced symbolic reasoning beyond GSM8k.
  • Guru (Cheng et al., 2025) is a multi-domain reasoning dataset with 91.9k examples spanning math, code, science, logic, simulation, and tabular tasks; we use its math subset (around 54k samples), which introduces diverse problem formats for evaluating transfer of reasoning strategies.
  • ToolACE (Liu et al., 2025a) is a multilingual benchmark with around 11k synthetic samples designed to evaluate LLMs’ ability to solve tasks by selecting and invoking external tools via strict JSON-formatted function calls; we use a 5k single-turn subset in our experiments.
@@ B.2 Understanding the synchronization parameters
The parameter sync_interval specifies the number of generated rollout batches (which equals the number of gradient steps for training the policy model) between two consecutive executions of model weight synchronization. When sync_interval =1=1, the rollout and policy models synchronize after each gradient step with one batch of samples, yielding a strictly on-policy process (if we ignore the issue of precision mismatch between rollout and training engines (Yao et al., 2025)). When sync_interval >1>1, sync_interval rollout batches are generated with stale model weights before synchronization, which accelerates the overall RL process through pipeline parallelism but incurs off-policyness.
The parameter sync_offset specifies the lag between the generation and consumption of each batch. More specifically, sync_offset batches are generated and saved to the buffer before training is launched, which is also useful for reducing pipeline bubbles and improving hardware utilization (Noukhovitch et al., 2025). In some of our experiments, we deliberately set sync_offset to a large value, in order to simulate a scenario where reward signals from the environment are lagged.
@@
GSM8K Qwen2.5 1.5B ToolACE Llama-3.2 3B Guru Qwen2.5 7B Guru Qwen3 30B-A3B MATH Llama-3.1 8B
Learning rate $\times\!10^{-6} 1×1$$\times\!10^{-6}$ $1\!\times\!10^{-6}$ $2\!\times\!10^{-6}$ $5\!\times\!10^{-7}$
Batch size 96 96 64 72 64
$K$ 8 8 16 16 16
Weight decay 0.01 0.01 0.1 0.1 0.1
Warmup steps 0 0 80 80 40
Eval temperature 1.0 N/A N/A N/A 0.6
Eval top-p 1.0 N/A N/A N/A 1.0
Figures 2, 4, 8, 9, 10 3 5 12 11
@@ B.3 REC with different clipping mechanisms
We replace the mask $M_{i}^{t}$ in REC-OneSide-IS/NoIS in Eq. (8) with a two-side mask [^3] It turns out that REC-TwoSide-NoIS resembles the sPPO algorithm proposed by Vaswani et al. (2022), though derived with different rationales. :
@@
Two-side clipping imposes weaker regularization than one-side clipping does with the same clipping parameter $(\epsilon_{\textsf{low}},\epsilon_{\textsf{high}})$. This can potentially improve training efficiency, but might also be risky when $\pi_{\bm{\theta}}/\pi_{\textsf{old}}$ goes far off. To compensate for this, we design REC-Ring.
@@ B.4 Ablation: the impact of learning rates
Remark 1.
Let us elaborate on the initial motivation and rationale behind our investigation of enlarging the clipping range. In the conventional theory of importance sampling and bias-variance trade-off for policy gradient estimation, we need certain trust-region condition that constrains the distance between the actor policy $\pi_{\bm{\theta}}$ and behavior policy $\pi_{\textsf{old}}$. By this rationale, LLM-RL frameworks commonly adopt small values $\epsilon_{\textsf{low}}=\epsilon_{\textsf{high}}=0.2$ for the default clipping range, or relax it up to $\epsilon_{\textsf{high}}=0.28$ as proposed by DAPO (Yu et al., 2025). Our work, in contrast, proposes a native off-policy interpretation for group-relative REINFORCE. The essential rationale is no longer about unbiased policy gradient estimation, and importance sampling is found to be non-essential. While clipping remains a valid option (among many others) for regularization, we are curious about whether the conventional choice of narrow clipping range is still necessary, especially after observing that GRPO’s learning progress slows down drastically (as the fraction of clipped tokens grows) with increasing off-policyness. Conceptually, our off-policy interpretation would advocate a larger clipping range than allowed by the conventional policy gradient theory. We hope that our preliminary exploration of this aspect could inspire the community to rethink the true working mechanism behind GRPO-like algorithms and the possibility of faster off-policy learning.
@@ B.5 Experiments for OPMD and AsymRE
Figure 10 presents empirical results for OPMD and AsymRE in various off-policy settings. It is worth noting that, while the analysis and experiments in their original papers (Kimi-Team, 2025b; Arnal et al., 2025) focus on a setting that is effectively the same as our $\texttt{sync\_interval}>1$ setting, our analysis and experiments have also validated their efficacy in $\texttt{sync\_offset}>0$ scenarios.
@@
B.6 Additional details and results for RED algorithms
We present further implementation details for the RED-Drop and RED-Weight algorithms investigated in Section 4.3:
  • RED-Weight: Each sample $i$ is weighted by $w_{i}=\exp({A_{i}}/{{T}})$, where $A_{i}$ denotes its advantage estimate and T>0{T}>0 is a temperature parameter controlling the sharpness of weighting. Intuitively, this scheme amplifies high-advantage samples while down-weighting low-advantage ones. We fix T=1{T}=1 for all experiments.
Additional experiments for RED-Weight, and its comparison against GRPO and REC-OneSide-NoIS, can be found in Figure 11. We observe that for the MATH dataset and Llama-3.1-8B-Instruct, RED-Weight achieves higher rewards with lower KL divergence, while maintaining more stable entropy and response lengths.
Figure 11: Comparison of RED-Weight, REC-OneSide-NoIS, and GRPO on MATH with Llama-3.1-8B-Instruct. Reported metrics for training include reward, KL divergence, entropy, and response length. We also report evaluation accuracy on the MATH500 subset.
Figure 12: Empirical results on Guru-Math with Qwen3-30B-A3B (MoE). Training reward curves are smoothed with a running-average window of size 3.
B.7 GSPO: sequence-level importance sampling could be non-essential
Group Sequence Policy Optimization (GSPO) (Zheng et al., 2025) proposes to replace token-wise clipping and importance sampling in GRPO with sequence-wise counterparts. Similar to Finding F1 in Section 4 for GRPO, we hypothesize that GSPO’s effectiveness stems from sequence-level clipping as regularization, rather than from sequence-level importance sampling. We provides preliminary validation for this hypothesis, through experiments with GSPO-style REC variants.
Implementations.
Given a prompt $x$ and $K$ responses $\{y_{i}\}_{1\leq i\leq K}$, let $s_{i}(\bm{\theta})$ denote the length-normalized sequence-level probability ratio for $y_{i}$:
$s_{i}(\bm{\theta})\coloneqq\bigg(\frac{\pi_{\bm{\theta}}(y_{i}\,|\,x)}{\pi_{\textsf{old}}(y_{i}\,|\,x)}\bigg)^{\frac{1}{|y_{i}|}}=\exp\bigg(\frac{1}{|y_{i}|}\sum_{1\leq t\leq|y_{i}|}\log\frac{\pi_{\bm{\theta}}(y_{i}^{t}\,|\,x,y_{i}^{<t})}{\pi_{\textsf{old}}(y_{i}^{t}\,|\,x,y_{i}^{<t})}\bigg).$
We further define the one-side sequence-level clipping mask as
$M_{i}\coloneqq\mathbbm{1}\bigg(A_{i}>0,\;s_{i}(\bm{\theta})\leq 1+\epsilon_{\textsf{high}}\bigg)+\mathbbm{1}\bigg(A_{i}<0,\;s_{i}(\bm{\theta})\geq 1-\epsilon_{\textsf{low}}\bigg).$
With these notations in place, we implement two GSPO-style REC variants as follows:
$\text{{REC-GSPO-IS}:}\;\;\bm{g}$ $=\frac{1}{K}\sum_{1\leq i\leq K}\frac{1}{|y_{i}|}\sum_{1\leq t\leq|y_{i}|}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}^{t}\,|\,x,y_{i}^{<t})\cdot(r_{i}-\overline{r})\,s_{i}(\bm{\theta})\,M_{i},$
$\text{{REC-GSPO-NoIS}:}\;\;\bm{g}$ $=\frac{1}{K}\sum_{1\leq i\leq K}\frac{1}{|y_{i}|}\sum_{1\leq t\leq|y_{i}|}\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}^{t}\,|\,x,y_{i}^{<t})\cdot(r_{i}-\overline{r})\,M_{i}.$
One can check that REC-GSPO-IS is equivalent to GSPO (except that we use $r_{i}-\overline{r}$ as the advantage, without normalization by $\sigma_{r}$), while REC-GSPO-NoIS discards the sequence-level importance-sampling weights.
Experiments.
We use the Guru-Math dataset and a mixture-of-expert (MoE) model — Qwen3-30B-A3B (Qwen-Team, 2025b) — since stable RL for MoE models is one of the main motivations behind GSPO (Zheng et al., 2025). We set sync_interval = 8, $\epsilon_{\textsf{low}}=6\times 10^{-4}$, and $\epsilon_{\textsf{high}}=8\times 10^{-4}$; other hyperparameters can be found in Tables 1 and 2.
Figure 12 shows that the learning curves of both REC-GSPO variants — with or without importance sampling — mostly overlap, indicating that importance sampling is likely a non-essential component for the effectiveness of GSPO.
@@ Appendix C Summary: a unified view of various algorithms
Add regularization loss OPMD $\widehat{L}=-\tfrac{1}{K}\!\sum_{i}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}|x)+\tfrac{{\beta}}{2K}\!\sum_{i}(\log\pi_{\bm{\theta}}(y_{i}|x)-\log\pi_{\textsf{old}}(y_{i}|x))^{2}$
AsymRE $\widehat{L}=-\tfrac{1}{K}\!\sum_{i}(r_{i}-\overline{r})\log\pi_{\bm{\theta}}(y_{i}|x)-\tfrac{{\beta}}{K}\!\sum_{i}\log\pi_{\bm{\theta}}(y_{i}|x)$
@@
RED-Weight $\bm{g}=\sum_{i}w_{i}(r_{i}-\overline{r})\,\nabla_{\bm{\theta}}\log\pi_{\bm{\theta}}(y_{i}|x),\;\;w_{i}=\exp(A_{i}/{T})$
Appendix D Additional related works
We focus our discussion on prior works that are most closely related to our core analysis in Section 2.2. In a tabular setting, the surrogate objective in Eq. (3) — KL-regularized reward maximization — can be regarded as an instantiation of mirror descent, whose optimum admits the closed form in Eq. (4). In more general settings with parameterized policy $\pi_{\bm{\theta}}$ and large action space, it is infeasible to realize Eq. (4) directly, and one would resort to optimizing the model parameters. Various algorithms have been developed on the basis of Eq. (3) and (4), including Kimi’s OPMD (Kimi-Team, 2025b) as explained in Section 4.2.
REBEL (Gao et al., 2024) has a derivation that largely overlaps with our Step 1 and 2 analysis in Section 2.2. It then seeks to solve the squared loss in Eq. (6), which enforces the pairwise consistency condition in Eq. (5). CoPG (Flet-Berliac et al., 2024) takes a similar approach, except that it uses a fixed reference policy (rather than the current iteration $\pi_{\bm{\theta}_{t}}$) for KL regularization. Compared to REINFORCE-style algorithms — for which enterprise‑grade LLM-RL frameworks like verl (Sheng et al., 2024) and Trinity-RFT (Pan et al., 2025) have been heavily optimized for — REBEL and CoPG could be less infrastructure-friendly or efficient. For example, in the presence of data parallelism and gradient accumulation, these frameworks can automatically divide a mini-batch into multiple micro-batches (each containing multiple or just one sequence) in a way that maximizes load balancing and training efficiency, while minimizing peak memory usage. However, solving the squared loss in Eq. (6) (like REBEL does) contradicts these performance optimization techniques, as it requires paired responses for the same prompt to be located within the same micro-batch. This constraint increases infrastructure complexity and peak memory usage, as reported by Brantley et al. (2025). Our Step 3 analysis in Section 2.2, on the other hand, proposes to take one gradient descent step for the squared loss, leading to a group-relative variant of classic REINFORCE while giving it a native off-policy interpretation.
Natural Policy Gradient (NPG) (Kakade, 2001) can be derived by approximating the surrogate objective in Eq. (3) with first-order Taylor expansion for the max-reward term and second-order Taylor expansion for the KL term, and then setting its gradient to zero. Since NPG requires on-policy sampling, it is less relevant to our study of off-policy LLM-RL. DPO (Rafailov et al., 2023) was also derived on the basis of Eq. (3), (4) and (5), but in a substantially different setting, with pairwise preference data and the Bradley-Terry assumption.